Determine the formula for the nth term of the sequence:




a) -2, 1, 7, 25, 79, ....




b) 0, 1, 3, 7, 15, ....


1
Expert's answer
2022-05-06T10:56:25-0400

a) Let "a_n=3^n-2, n\ge0, n\in \Z." Then


"a_0=-1, a_1=1, a_2=7, a_3=25,a_4=79,..."

"a_n=3^n-2, n\ge0, n\in \Z"


b) Let "a_n=2^n-1, n\ge0, n\in \Z." Then


"a_0=0, a_1=1, a_2=3, a_3=7,a_4=15,..."

"a_n=2^n-1, n\ge0, n\in \Z"